100.0% complete

This is a Single Choice Question; skip ahead to question content

A

B

C

D

Confirm

In Year 1, Lena invests $80,000 for a 20% partnership interest in an activity in which she works full time. The partnership reports losses of $300,000 in Year 1 and $150,000 in Year 2. Lena’s share of the partnership’s losses is $60,000 in Year 1 and $30,000 in Year 2. How much, if any, of Lena’s losses will be suspended?

  1. No suspended losses; no at risk losses.
  2. Year 1 passive losses suspended $60,000; Year 2 $20,000 passive loss suspended and $10,000 at risk suspended.
  3. Year 1 no suspended losses; Year 2 $30,000 suspended for at risk.
  4. Year 1 no suspended losses; Year 2 $20,000 passive loss and $10,000 suspended for at risk.
Answer #1

The answer is A – No suspended losses, no at risk losses.

REASON:

Sec. 704(d) provides that a partner’s distributive share of loss is allowable to the extent of the partner’s adjusted tax basis in his interest in the partnership at the end of the partnership year in which the loss occurred. Any losses in excess of the partner’s tax basis are disallowed pro rata (Regs. Sec. 1.704-1(d)) and are carried forward indefinitely for as long as the partner remains in the partnership.

Thus, in the given question-

               Share of Lena = 20%

               Share in loss of 1st year = 300,000*20% = 60,000

               Share in loss of 2st year = 150,000*20% = 30,000

Also, the allocated losses are the same, i.e., 60,000 and 30,000.

This concludes that there are no suspended losses at all as:

                              Loss allocated = applicable loss to be allocated

Moreover, there are no at risk losses as the information given in the question doesn’t satisfy any of the following conditions:

  • the amount of money and the adjusted basis of property contributed to an activity;
  • amounts borrowed with respect to the activity to the extent the taxpayer is personally liable for repayment or has pledged property, other than property used in the activity, as security for the borrowed amount; and
  • generally, amounts borrowed with respect to the activity of holding real property for which no person is personally liable for repayment (qualified nonrecourse financing).